• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ransome

    Log entry time 15:51:03 on June 1,2002

    Entry number 4385

    keyword=shift summary

    8:00 begin shift with dipoles being reset
    Left dipole tripped just before 8:00
    We let run down to zero and turned on.
    Came up part way, then stopped and no longer responds.
    Also, NMR doesn't lock or respond
    Right dipole comes up short (.98980 instead of .9922), then stopped moving
    Also does not respond to changes.

    8:40 Called Mark Stevens for help
    9:20 Mark Stevens arrives
    10:30 out of the hall

    Right arm had lost all communication with dipole, had to
    turn off and recycle magnet.

    Left arm had lost communication on dipole and 2 quads.
    All turned off and recycled (including high current cycle of Q2 and Q3).

    10:45 beam back at 5 uA, but not yet centered on BPM's
    We prepare to do harp scan target is thin 12C
    NCD Control screen dead
    Also had to reboot target control IOC, now OK
    11:20 Harp scan done run 1762, CODA run 1053
    Found size too small, MCC trying to adust beam
    Run 1763, now beam is too large
    12:00 beam off due to RF problems
    12:20 beam back
    12:40 we now have raster set for 3 x 6 (which gives about 2 x 2 on spot)
    12:44 ask for 100 uA on thin 12C
    13:15 checked detectors with 12C - looks OK, go to LD2 15 cm
    Begin current studies -
    Lowest current MCC can do is 2.7 uA, we start there
    We are also skipping runs at 4, and 6.5 uA.
    Will also take 100 Hz of trigger T5
    14:53 beam off RF problems (run 1066)
    15:07 beam back
    CODA crash

    Run number current
    1058 2.8 uA
    1059 5.0 uA
    1060 8.7 uA (beam off briefly after 450,000 events)
    1061 11.0 uA
    1062 15.8 uA (stopped, dead time too high)
    1063 15.8 uA
    1064 19.9 uA
    1065 29.5 uA
    1066 39.2 uA (beam off, stopped after few events)
    1067 39.2 uA
    1068 49.5 uA
    1069 69.5 uA (stopped, lost contact with all ROC, TS)